# Architecting mechanisms of damage in topological metamaterials

### v1

@@ Abstract
Architecting mechanisms of damage in metamaterials by leveraging lattice topology and geometry poses a vital yet complex challenge, essential for engineering desirable mechanical responses. Of these metamaterials, Maxwell lattices, which are on the verge of mechanical stability, offer significant potential for advanced functionality. By leveraging their robust topological features, they enable precise control of effective elastic properties, manipulation of stress localisation and delocalisation across specific domains, and targeted global damage that follows local fracture events. In this work, we identify topology and geometry-dependent parameters that establish a simple, yet precise, framework for designing the behavior of non-idealized Maxwell lattices and their damage processes. Using a beam element simulator, we systematically explore the underlying phenomenology to demonstrate how this framework can guide or arrest damage in lattices, both with and without domain walls and additional boundary constraints. Our approach uncovers a robust way to manipulate the mechanisms of damage and the path they follow in metamaterials, with further insight into crack arrest, diversion, and shielding.
Mechanical metamaterial |||| Topological mechanics |||| Topological polarisation |||| Damage |||| Fracture
preprint: APS/123-QED
Introduction
@@
Significant efforts have been made to determine the relationship between fracture toughness and the underlying microstructure of lattice-based materials [1, 2, 3, 4, 5, 6, 7, 8]. Beyond metrics related to the onset of failure, the resistance to crack growth during the damage propagation has also been explored [9, 10, 11, 12], highlighting an opportunity to increase their steady-state fracture toughness. Harnessing the change in failure characteristics of materials straddling the rigidity transition [13], where the fracture delocalises, may also provide a pathway to increase this toughening behaviour. Attempts to design the damage evolution have generally focused on the inclusion of heterogeneities to control the percolating pathway through: material property [14] or density [15] variations, anisotropic unit cell design [16, 17], and disorder [18]. The focus on leveraging heterogeneities to influence damage mechanisms have over-passed the rich design space within their homogeneous counterparts. As a result, limited resources are available to accurately predict the stress profile and damage evolution in an already damaged homogeneous lattice-based material, in terms of simplified intuitive principles, derived from the underlying microstructure.
@@
A general indication of the dominant behaviour of a lattice is given by its coordination number, $z$. In d dimensions, lattices with $z<2d$ ($z>2d$), are bending (stretching)-dominated [19]. Maxwell Lattices are positioned between these two extremes, on the verge of mechanical stability, containing an equal number of nontrivial zero energy modes, or floppy modes (FM), and states of self-stress (SSS). By bridging the gap between Maxwell Lattices and the topological band theory of electronic systems, Kane and Lubensky [20] showed that the localisation of these surface FM and SSS can be manipulated through a type of topological polarisation $\textbf{P}_{\mathrm{T}}$, that depends on the unit-cell’s geometry. The robustness of this topological quantity provides a powerful tool that can be leveraged to architect the deformation and stress localisation in these lattices [21, 13, 22].
@@
Building upon concepts derived from the robustness of topological properties, efforts to control the mechanics of damage processes in Maxwell lattices have yielded significant results, mainly through the introduction of SSS achieved by combining topologically distinct domains [23, 24, 25, 26]. For instance, in idealised (zero-thickness hinges) lattices, it was shown through numerical simulations that SSS domain walls can be used to focus the stress and prevent failure from initiating at imperfections [23]. However, when these lattices are manufactured in practice, the joints will inevitably exhibit a non-zero bending rigidity (non-idealised). This was shown to dilute the stress focusing effect [24, 25, 26] through the delocalisation of the stress associated with next nearest neighbour (NNN) coupling, analogous to phenomena described in the case of brittle materials [27]. By considering NNN interactions, it has been illustrated how the unit-cell geometry can be tuned to adjust this delocalisation [24], although the damage evolution has not yet been investigated. Furthermore, while the inclusion of NNN provides insights into the delocalisation effects, further discretisation is required to accurately capture the behaviour and damage propagation, particularly at larger values of the unit-cell size to strut thickness, i.e. slenderness ratio (SR), where bending energy becomes an important contributor.
@@
As damage propagates through a lattice, constraints are removed as the individual elements fracture, necessarily resulting in the removal of SSS and an emergence of FM, thus establishing a certain compliance in the system. Understanding how to control the form of these additional modes opens up a vast design space that amends to the role of topology. Although the edge localization is topologically robust, there is great latent potential for adjusting the form of these modes in lattices that fall under the same equivalence class, which is defined uniquely through the direction of the polarisation vector $\textbf{P}_{\mathrm{T}}$. This, in turn, provides a yet unrealized tool for precisely controlling the stress concentration and thus a mechanism that can guide the damage path, without necessarily altering $\textbf{P}_{\mathrm{T}}$. Furthermore, while it has been shown that SSS domain walls can focus stress [23, 24, 25, 26], to the best of our knowledge no insights have yet been provided on how FM, due to imperfections, influence the development and damage of lattices with or without these SSS domain walls.
@@
Background Theory
The Role of Topology
The global count of FM ($N_{\mathrm{0}}$) and SSS ($N_{\mathrm{s}}$) in a lattice are quantities determined by the topology of the network itself, and are captured by the Maxwell-Calladine index theorem [28, 29], which relates them to the number of sites ($N$) and constraints ($N_{\mathrm{c}}$) in the system:
@@
To that end, a kinematic analysis of the unit-cell can be used to connect the site displacements, $\mathbf{u}$, to the bond elongations, $\mathbf{e}$, through the compatibility matrix $\mathbf{C}$ (i.e. $\mathbf{e}=\mathbf{C}\cdot\mathbf{u}$). The FM can then be found in the nullspace of $\mathbf{C}$, signifying combinations of node displacements that do not stretch any bonds. Likewise, the site forces, $\mathbf{f}$, can be related to the bond tensions, $\mathbf{t}$, using the equilibrium matrix $\mathbf{Q}$ (i.e. $\mathbf{f}=\mathbf{Q}\cdot\mathbf{t}$). Therefore, the nullspace of $\mathbf{Q}$ contains the SSS. Invoking the principle of virtual work, it can be shown that $\mathbf{C}=\mathbf{Q}^{\mathrm{\dagger}}$ [29], implying that FM and SSS form conjugate pairs. Moreover, depending on the approximation within which the system is analysed, different forms for these modes can be unveiled. For example, a Bloch wave expansion of the lattice distortions can capture all possible periodic modes and their energetic costs, with the latter contained in the lattices dispersion relation [30]. There is also a certain class of affine strain producing mechanisms, commonly referred to as Guest-Hutchinson (GH) modes [19], which can be unveiled through augmenting C to allow for boundary deformations of the unit cell.
From (1), Maxwell lattices can be defined in periodic boundary conditions as having an equal number of constraints and degrees of freedom, resulting in $N_{\mathrm{0}}=N_{\mathrm{s}}$. In practice, constructing a finite specimen from an infinite periodic domain will necessarily change the number of constraints, and the specimen will lose/gain FM/SSS, depending on the boundary conditions. For example, Maxwell Lattice specimens with free boundaries and a gapped dispersion relation, will have to possess extra evanescent FM, i.e. modes that exponentially localise on bulk discontinuities, such as free surfaces [20]. Importantly, these edge modes can exhibit a topologically protected directional preference, also known as a topological polarization, asymmetrically changing the lattice response depending on which surface they localise.
@@
where $\textbf{r}_{\mathrm{\mu}}^{\mathrm{0}}$ are the basis vectors for the sites of the RK, $\textbf{p}_{\mathrm{\mu}}$ are the vectors normal to the bonds, $\textbf{a}_{\mathrm{\mu}}$ are the primitive lattice vectors, and $\mu\in\{1,2,3\}$ [31, 32]. This parameterisation displaces the sites so that each of the three originally straight filaments of bonds in RK is disrupted individually by each one of the $x_{\mu}$, while $z_{\mathrm{s}}$ allows for the scaling of neighbouring triangles. A description of the lattice using the unit-cell defined by (2), allows the topological polarization to be fully captured by a simple expression only depending on the sign of $x_{\mu}$ [20] when $\mathbf{a}_{\mathrm{\mu}}$ is fixed:
@@
Our simulations are focused on finite kagome lattices with a pre-crack located along the mid-plane of the lattice, parallel to the x-axis. Creating the pre-crack amounts to removing two constraints per unit cell, which, depending on the boundary conditions, will either introduce two FM localising on the crack edges or remove two SSS from the top/bottom boundaries of the system. Fixing $\mathbf{a}_{\mathrm{\mu}}$ and noting the mirror symmetry of the domain, there are only 6 non-redundant triplets of $\mathrm{sgn}(x_{\mathrm{1}},x_{\mathrm{2}},x_{\mathrm{3}})$, differentiating the relevant equivalence classes through $\textbf{P}_{\mathrm{T}}$. Namely (+,+,+)(+,+,+)( + , + , + ) or (,,)(-,-,-)( - , - , - ), (,+,+)(-,+,+)( - , + , + ), (+,,)(+,-,-)( + , - , - ), (,+,)(-,+,-)( - , + , - ) and (+,,+)(+,-,+)( + , - , + ) which we refer to as left (LTK) or right (RTK) twisted kagome, left polarised (LP), right polarised (RP), left-up polarised (LUP), and right-down polarised (RDP). We also consider the RK lattice, which represents a transition point between equivalence classes where the topological index is not defined.
The Role of Geometry
@@
Utilising the zero-energy condition for the non-trivial FM (SSS), or equivalently observing that they fall in the nullspace of $\mathbf{C}$ ($\mathbf{Q}$), restricts their possible functional form. This constraint, which can be expressed in Fourier Space as $E=\det\mathbf{C}(q_{\textrm{1}},q_{\textrm{2}})=\det\mathbf{Q}(q_{\textrm{1}},q_{\textrm{2}})=0$, amounts to fixing one wavenumber component in terms of the other, effectively reducing the degrees of freedom that are relevant for their description. This effect, known as bulk-boundary correspondence, allows the reconstruction of the mode in the bulk given knowledge of its amplitude on a boundary. In essence, an evanescent mode that is described by a real wavenumber $q_{\mathrm{1}}$ parallel to a boundary will be associated with a wavenumber $q_{\mathrm{2}}\equiv q_{\mathrm{2}}(q_{\mathrm{1}})$ orthogonal to it [20]. If the dispersion relation is gapped, the solution for the orthogonal component will be complex, $q_{\mathrm{2}}=k_{\mathrm{2}}+i\,\kappa_{\mathrm{2}}$, where $k_{\mathrm{2}}$ is the real part of the wavenumber and $\kappa_{\mathrm{2}}$, its imaginary part, is the inverse decay length in the $2$-direction. Therefore, a sinusoidal disturbance decays into the bulk with respect to the edge’s inward facing normal at an angle given by
@@
with clockwise rotation for a negative angle. As $\mathbf{C}=\mathbf{Q}^{\dagger}$, the $q_{\mathrm{2}}$ component of the SSS will be the complex conjugate of the FM resulting in $\kappa^{\mathrm{SSS}}=-\kappa^{\mathrm{FM}}$ and $\theta^{\mathrm{SSS}}=\theta^{\mathrm{FM}}=\theta$.
@@
Finally, the theory summarised in this section is derived assuming lattices constructed from springs that are free to rotate at their ends. Practically, it is difficult to construct a lattice with these restrictions, and in reality, non-zero bending resistance will occur. This bending resistance penalises the FM and SSS, resulting in these modes lifting from zero energy, weakening the effect of the polarization [24] through the delocalisation effect. As per the definition of the slenderness ratio, $\mathrm{SR}=|\mathbf{a}_{\mathrm{1}}|/t$, where $t$ is the bond thickness and $|\mathbf{a}_{\mathrm{1}}|$ is the magnitude of the primitive lattice vector in the 1-direction, this quantity is identified as an additional geometrical parameter that controls the global behavior.
Behaviour of Topologically Distinct Lattices

### v2

@@ Abstract
Architecting mechanisms of damage in metamaterials by leveraging lattice topology and geometry poses a vital yet complex challenge, essential for engineering desirable mechanical responses. Of these metamaterials, Maxwell lattices, which are on the verge of mechanical stability, offer significant potential for advanced functionality. By leveraging their robust topological features, they enable precise control of effective elastic properties, manipulation of stress localisation and delocalisation across specific domains, and targeted global damage that follows local fracture events. In this work, we identify topology and geometry-dependent parameters that establish a simple, yet precise, framework for designing the behaviour of non-idealised Maxwell lattices and their damage processes. We numerically explore the underlying phenomenology to demonstrate how this framework can guide or arrest damage in lattices, both with and without domain walls and additional boundary constraints. Our approach uncovers a robust way to manipulate the mechanisms of damage and the path they follow in metamaterials, with further insight into crack arrest, diversion, and shielding.
keywords:
Mechanical metamaterial , Topological mechanics |||| Topological polarisation , Damage , Fracture
[…]
organization=Institute for Infrastructure and Environment, School of Engineering, University of Edinburgh,addressline=, city=Edinburgh, postcode=EH9 3FG, state=, country=UK
1 Introduction
@@
Significant efforts have been made to determine the relationship between fracture toughness and the underlying microstructure of lattice-based materials (Gibson, 2003; Chen et al., 1998; Fleck and Qiu, 2007; Tankasala et al., 2015; Athanasiadis et al., 2021; Berkache et al., 2022; Omidi and St-Pierre, 2023; Choukir and Singh, 2023). Beyond metrics related to the onset of failure, the resistance to crack growth during the damage propagation has also been explored (Schmidt and Fleck, 2001; Tankasala and Fleck, 2020; Hsieh et al., 2020; Wang et al., 2024), highlighting an opportunity to increase their steady-state fracture toughness. Harnessing the change in failure characteristics of materials straddling the rigidity transition (Driscoll et al., 2016), where the fracture delocalises, may also provide a pathway to increase this toughening behaviour. Attempts to design the damage evolution have generally focused on the inclusion of heterogeneities to control the percolating pathway through: material property (Gao et al., 2020) or density (Manno et al., 2019) variations, anisotropic unit cell design (Domino et al., 2024; Gao et al., 2024), and disorder (Karapiperis and Kochmann, 2023). The focus on leveraging heterogeneities to influence damage mechanisms have over-passed the rich design space within their homogeneous counterparts. As a result, limited resources are available to accurately predict the stress profile and damage evolution in an already damaged homogeneous lattice-based material, in terms of simplified intuitive principles, derived from the underlying microstructure.
@@
A general indication of the dominant behaviour of a lattice is given by its coordination number, $z$. In d dimensions, lattices with $z<2d$ ($z>2d$), are bending (stretching)-dominated (Guest and Hutchinson, 2003). Maxwell Lattices are positioned between these two extremes, on the verge of mechanical stability, containing an equal number of nontrivial zero energy modes, or floppy modes (FM), and states of self-stress (SSS). By bridging the gap between Maxwell Lattices and the topological band theory of electronic systems, (Kane and Lubensky, 2014) showed that the localisation of these surface FM and SSS can be manipulated through a type of topological polarisation $\textbf{P}_{\mathrm{T}}$, that depends on the unit-cell’s geometry. The robustness of this topological quantity provides a powerful tool that can be leveraged to architect the deformation and stress localisation in these lattices (Paulose et al., 2015; Driscoll et al., 2016; Chapuis et al., 2022).
@@
Building upon concepts derived from the robustness of topological properties, efforts to control the mechanics of damage processes in Maxwell lattices have yielded significant results, mainly through the introduction of SSS achieved by combining topologically distinct domains (Zhang and Mao, 2018; Liu et al., 2023; Widstrand et al., 2023, 2024). For instance, in idealised (zero-thickness hinges) lattices, it was shown through numerical simulations that SSS domain walls can be used to focus the stress and prevent failure from initiating at imperfections (Zhang and Mao, 2018). However, when these lattices are manufactured in practice, the joints will inevitably exhibit a non-zero bending rigidity (non-idealised). This was shown to dilute the stress focusing effect (Liu et al., 2023; Widstrand et al., 2023, 2024) through the delocalisation of the stress associated with next nearest neighbour (NNN) coupling, analogous to phenomena described in the case of brittle materials (Salman and Truskinovsky, 2021). By considering NNN interactions, it has been illustrated how the unit-cell geometry can be tuned to adjust this delocalisation (Liu et al., 2023), although the damage evolution has not yet been investigated. Furthermore, while the inclusion of NNN provides insights into the delocalisation effects, further discretisation is required to accurately capture the behaviour and damage propagation, particularly at larger values of the unit-cell size to strut thickness, i.e. slenderness ratio (SR), where bending energy becomes an important contributor.
@@
As damage propagates through a lattice, constraints are removed as the individual elements fracture, necessarily resulting in the removal of SSS and an emergence of FM, thus establishing a certain compliance in the system. Understanding how to control the form of these additional modes opens up a vast design space that amends to the role of topology. Although the edge localization is topologically robust, there is great latent potential for adjusting the form of these modes in lattices that fall under the same equivalence class, which is defined uniquely through the direction of the polarisation vector $\textbf{P}_{\mathrm{T}}$. This, in turn, provides a yet unrealised tool for precisely controlling the stress concentration and thus a mechanism that can guide the damage path, without necessarily altering $\textbf{P}_{\mathrm{T}}$. Furthermore, while it has been shown that SSS domain walls can focus stress (Zhang and Mao, 2018; Liu et al., 2023; Widstrand et al., 2023, 2024), to the best of our knowledge no insights have yet been provided on how FM, due to imperfections, influence the development and damage of lattices with or without these SSS domain walls.
@@
2 Background Theory
2.1 The Role of Topology
The global count of FM ($N_{\mathrm{0}}$) and SSS ($N_{\mathrm{s}}$) in a lattice are quantities determined by the topology of the network itself, and are captured by the Maxwell-Calladine index theorem (Maxwell, 1864; Calladine, 1978), which relates them to the number of sites ($N$) and constraints ($N_{\mathrm{c}}$) in the system:
@@ 2.1 The Role of Topology
To that end, a kinematic analysis of the unit-cell can be used to connect the site displacements, $\mathbf{u}$, to the bond elongations, $\mathbf{e}$, through the compatibility matrix $\mathbf{C}$ (i.e. $\mathbf{e}=\mathbf{C}\cdot\mathbf{u}$). The FM can then be found in the nullspace of $\mathbf{C}$, signifying combinations of node displacements that do not stretch any bonds. Likewise, the site forces, $\mathbf{f}$, can be related to the bond tensions, $\mathbf{t}$, using the equilibrium matrix $\mathbf{Q}$ (i.e. $\mathbf{f}=\mathbf{Q}\cdot\mathbf{t}$). Therefore, the nullspace of $\mathbf{Q}$ contains the SSS. Invoking the principle of virtual work, it can be shown that $\mathbf{C}=\mathbf{Q}^{\mathrm{\dagger}}$ (Calladine, 1978), implying that FM and SSS form conjugate pairs. Moreover, depending on the approximation within which the system is analysed, different forms for these modes can be unveiled. For example, a Bloch wave expansion of the lattice distortions can capture all possible periodic modes and their energetic costs, with the latter contained in the lattices dispersion relation (Hutchinson and Fleck, 2006). There is also a certain class of affine strain producing mechanisms, commonly referred to as Guest-Hutchinson (GH) modes (Guest and Hutchinson, 2003), which can be unveiled through augmenting C to allow for boundary deformations of the unit cell.
From (1), Maxwell lattices can be defined in periodic boundary conditions as having an equal number of constraints and degrees of freedom, resulting in $N_{\mathrm{0}}=N_{\mathrm{s}}$. In practice, constructing a finite specimen from an infinite periodic domain will necessarily change the number of constraints, and the specimen will lose/gain FM/SSS, depending on the boundary conditions. For example, Maxwell Lattice specimens with free boundaries and a gapped dispersion relation, will have to possess extra evanescent FM, i.e. modes that exponentially localise on bulk discontinuities, such as free surfaces (Kane and Lubensky, 2014). Importantly, these edge modes can exhibit a topologically protected directional preference, also known as a topological polarization, asymmetrically changing the lattice response depending on which surface they localise.
@@
where $\textbf{r}_{\mathrm{\mu}}^{\mathrm{0}}$ are the basis vectors for the sites of the RK, $\textbf{p}_{\mathrm{\mu}}$ are the vectors normal to the bonds, $\textbf{a}_{\mathrm{\mu}}$ are the primitive lattice vectors, and $\mu\in\{1,2,3\}$ (Mao and Lubensky, 2018; Lubensky et al., 2015). This parameterisation displaces the sites so that each of the three originally straight filaments of bonds in RK is disrupted individually by each one of the $x_{\mu}$, while $z_{\mathrm{s}}$ allows for the scaling of neighbouring triangles. A description of the lattice using the unit-cell defined by (2), allows the topological polarization to be fully captured by a simple expression only depending on the sign of $x_{\mu}$ (Kane and Lubensky, 2014) when $\mathbf{a}_{\mathrm{\mu}}$ is fixed:
@@
Our simulations are focused on finite kagome lattices with a pre-crack located along the mid-plane of the lattice, parallel to the x-axis. Creating the pre-crack amounts to removing two constraints per unit cell, which, depending on the boundary conditions, will either introduce two FM localising on the crack edges or remove two SSS from the top/bottom boundaries of the system. Fixing $\mathbf{a}_{\mathrm{\mu}}$ and noting the mirror symmetry of the domain, there are only 6 non-redundant triplets of $\mathrm{sgn}(x_{\mathrm{1}},x_{\mathrm{2}},x_{\mathrm{3}})$, differentiating the relevant equivalence classes through $\textbf{P}_{\mathrm{T}}$. Namely (+,+,+) or (-,-,-), (-,+,+), (+,-,-), (-,+,-) and (+,-,+) which we refer to as left (LTK) or right (RTK) twisted kagome, left polarised (LP), right polarised (RP), left-up polarised (LUP), and right-down polarised (RDP). We also consider the RK lattice, which represents a transition point between equivalence classes where the topological index is not defined.
2.2 The Role of Geometry
@@
Utilising the zero-energy condition for the non-trivial FM (SSS), or equivalently observing that they fall in the nullspace of $\mathbf{C}$ ($\mathbf{Q}$), restricts their possible functional form. This constraint can be expressed in Fourier Space as
$E(q_{\mathrm{1}},q_{\mathrm{2}})=\det\mathbf{C}(q_{\textrm{1}},q_{\textrm{2}})=\det\mathbf{Q}(q_{\textrm{1}},q_{\textrm{2}})=0,$ (4)
which amounts to fixing one wavenumber component in terms of the other, effectively reducing the degrees of freedom that are relevant for their description. This effect, known as bulk-boundary correspondence, allows the reconstruction of the mode in the bulk given knowledge of its amplitude on a boundary (Fig. 1). In essence, an evanescent mode that is described by a real wavenumber $q_{\mathrm{1}}$ parallel to a boundary will be associated with a wavenumber $q_{\mathrm{2}}\equiv q_{\mathrm{2}}(q_{\mathrm{1}})$ orthogonal to it (Kane and Lubensky, 2014). If the dispersion relation is gapped, the solution for the orthogonal component will be complex, $q_{\mathrm{2}}=k_{\mathrm{2}}+i\,\kappa_{\mathrm{2}}$, where $k_{\mathrm{2}}$ is the real part of the wavenumber and $\kappa_{\mathrm{2}}$, its imaginary part, is the inverse decay length in the $2$-direction. Therefore, a sinusoidal disturbance decays into the bulk with respect to the edge’s inward facing normal at an angle given by
@@
with clockwise rotation for a negative angle (Fig. 1). As $\mathbf{C}=\mathbf{Q}^{\dagger}$, the $q_{\mathrm{2}}$ component of the SSS will be the complex conjugate of the FM resulting in $\kappa^{\mathrm{SSS}}=-\kappa^{\mathrm{FM}}$ and $\theta^{\mathrm{SSS}}=\theta^{\mathrm{FM}}=\theta$.
@@
Finally, the theory summarised in this section is derived assuming lattices constructed from springs that are free to rotate at their ends. Practically, it is difficult to construct a lattice with these restrictions, and in reality, non-zero bending resistance will occur. This bending resistance penalises the FM and SSS, resulting in these modes lifting from zero energy, weakening the effect of the polarization (Liu et al., 2023) through the delocalisation effect. As per the definition of the slenderness ratio, $\mathrm{SR}=|\mathbf{a}_{\mathrm{1}}|/t$, where $t$ is the bond thickness and $|\mathbf{a}_{\mathrm{1}}|$ is the magnitude of the primitive lattice vector in the 1-direction, this quantity is identified as an additional geometrical parameter that controls the global behaviour.
Figure 1: Discrete and continuum surface (floppy) mode visualizations for LP1. Green solid lines represent free edges, with a distortion on the edge made up of a single wavenumber $q_{\mathrm{1}}=2\pi s/N$, with a visualization of the distorted lattice fading into the continuum. The continuum plot depict the modulus $|\phi|$, of the otherwise complex FM amplitude (details of the continuum visualisation can be found in B).
3 Numerical Method
In the simulations, we represent the lattice as a mass-spring network with both axial and torsional springs providing NN and NNN interactions (Fig. 2a. The energy in this system is defined by
where: $N_{\mathrm{b}}$ and $N_{\mathrm{r}}$ are the number of axial and torsional springs in the system; $k_{\mathrm{i}}$ and $c_{\mathrm{s}}$ are the axial and torsional spring stiffness (defined in A.2); $l_{\mathrm{i}}$ and $l_{\mathrm{i}}^{\mathrm{0}}$ are the current length and original length of the axial spring; and, $\alpha_{\mathrm{s}}$ and $\alpha_{\mathrm{s}}^{\mathrm{0}}$ are the current angle and original angle between bonds $i$ and $l$. Notice that this formulation captures large deformations. $l_{\mathrm{i}}$ and $\alpha_{\mathrm{s}}$ can be written in terms of the site degrees of freedom $x_{\mathrm{j}}$ and, within each applied displacement step $\delta$, we minimises the force at each site ($\partial E/\partial x_{\mathrm{j}}$) below a specified tolerance ($F_{\mathrm{tol}}$) using the FIRE 2.0 algorithm (Bitzek et al., 2006; Guénolé et al., 2020).
To accurately capture the bending of the bonds, we further discretise the bonds into three elements (Fig. 2b). This level of discretisation was found to sufficiently capture the response of the lattice for the SR and material properties simulated.
Figure 2: (a) NN and NNN bonds. (b) Discretisation of the lattice.
An element has reached failure if
where $\sigma_{\mathrm{A}}$ and $\sigma_{\mathrm{B}}$ are the axial and bending stresses in the element (defined in A.3), and $\sigma_{\mathrm{A}}^{\mathrm{u}}$ and $\sigma_{\mathrm{B}}^{\mathrm{u}}$ are the material failure stress values in axial and bending respectively. While only the material failure is included explicitly, buckling related failure modes are captured by this model. When a bond becomes unstable, further rotation is possible without significantly increasing the applied force (El Nashie, 1990). This far-from-threshold rotation will eventually result in the failure of an element, which is simulated by splitting the site associated with the failed element into two, creating a new edge within the lattice. When either side of the bond meets the failure criteria, we separate the bond(s) by splitting the site into two and reassigning the bonds to this new site. This is done arbitrarily when two elements are connected to a site. The associated torsional spring stiffness ($c_{\mathrm{s}}$) is also set to 0. In Fig. 3a we show this process, where a yellow circle represents the new site. When three or four elements are connected at a site we choose to disconnect the elements about the two torsional springs whose angle changes the most i.e. if two elements are rotating together about the site they will separate from the other element(s) together (Fig. 3b,c). $c_{\mathrm{s}}$ of the two torsional springs associated with the largest rotations is then also set to 0. If, in this step, one element is separated from three elements, a new torsional spring is added to the set of three elements, connecting the adjacent elements that were previously on either side of the single separated element (Fig. 3c). No torsional springs are added to the system if only two elements remain after splitting the site (Fig. 3b,c). Further information about the method can be found in A.
Figure 3: Splitting of sites to simulate the fracture process. Splitting of sites with: (a) two NN bonds; (b) three NN bonds; and, (c) four NN bonds.
The material properties are taken from that of a relatively rigid, but brittle, material (Luan et al., 2022), with $E=10$GPa, $\sigma_{\mathrm{A}}^{\mathrm{u}}=65$MPa and $\sigma_{\mathrm{B}}^{\mathrm{u}}=126$MPa. The maximum allowable residual force ($F_{\mathrm{tol}}$) was set to 0.0025N in the simulations following a sensitivity study.
4 Behaviour of Topologically Distinct Lattices